Activity based costing and management

Opsonin Inc manufactures three products for the pharmaceuticals industry . product P annual sales, 8000 units. product Q annual sales, 15 000 units. product R : annual sales, 4000 units. The company uses a traditional, volume-based product costing system with manufacturing overhead applied on the basis of direct labour dollars. The product costs have been calculated as follows: Product P Product Q Product R $52.50 12.00 (0.6 hr x $20) 105.00 ($12x875%) Raw material Direct labour Manufacturing overhead* Total product cost S35.00 16.00 (0.8 hrx S20) 140.00 ($ 16x875%) $17.50 8.00 (0.4 hr x S20) 70.00 ($8x875%) $95.50 $169.50 S191.00 *Manufacturing overhead budget: Machinery Machine setup Inspection Material handling Engineering Total $1 225 000 5 250 525 000 875 000 2 975 000 Opsonin Inc's pricing method has been to set a budgeted setting price equal to 150 per cent of full product cost. However, only the Product Q have been selling at their budgeted price. The budgeted and actual current prices for all three products are the following: Product P $191.00 286.50 213.00 Product Q $169.50 254.25 254.25 Product R $ 95.50 143.25 200.00 Product cost Budgeted price Actual current selling price Opsonin Inc has been forced to lower the price of product P in order to get orders. In contrast, Opsonin Inc has raised the price of product R several times, but there has been no apparent loss of sales. Opsonin Inc has been under increasing pressure to reduce the price even further on gismos. In contrast, Opsonin Inc's competitors do not seem to be interested in the market for Product R. Opsonin Inc apparently has this market to itself. Required: 1. Is product P the company's least profitable product? Explain your answer 2. Is product R a profitable product for Opsonin Ltd? Explain your answer 3. Comment on the reactions of Opsonin Inc's competitors to the firm's pricing strategy. What 1 Marks] 1 Marks 2 Marks dangers does Opsonin Inc face?
4. Opsonin Inc's financial controller, Obrien Sally, recently attended a conference at which activity-based costing systems were discussed. She became convinced that such a system would help Opsonin Inc's management to understand its product costs better. She obtained top management's approval to design an activity-based costing system, and an ABC project team was formed. In Stage 1 of the ABC project, each of the overhead items listed in the overhead budget was placed into its own activity cost pool. Then an activity driver was identified for each activity cost pool. Finally, the ABC project team compiled data showing the percentage of each activity driver that was consumed by each of Opsonin Inc's product lines. These data are summarised as follows: Activity cost pool Activity driver Machine Machine setup Inspection Material handling Raw material costs Engineering product P 25% 20% 15% 25% 35% roduct Q 50% 30% 45% 69% 10% product R 25% 50% 40% Machine hours Number of setups Number of inspections ry 6% 55% Number of change orders Show how the financial controller determined the percentages given above for raw material [4 Marks] 5. Develop product costs for the three products on the basis of a simple activity-based product 3 Marks] 6. Calculate a budgeted price for each product, using Opsonin Inc's pricing formula. Compare the new budgeted prices with the current actual selling prices and previously reported 2 Marks] 7. Refer to the new budgeted prices for Opsonin Inc's three products, based on the new activity- based costing system. Write a memo to the company managing director commenting on the situation Opsonin Inc has been facing regarding the market for its products and the actions of its competitors. Discuss the strategic options available to management. What do you costs. (Round to the nearest whole per cent.) costing system. (Round to the nearest cent.) product costs recommend, and why? 5 Marks 8. Refer to the product costs developed in requirement 5. Prepare a table showing how Opsonin Inc's traditional, volume-based product costing system distorts the product costs of product 2 Marks] P, Q and R
Answer #1
1 As per traditional overhead allocation ,P is the Least profitable product
If we consider direct costsonly, P is more profitable than Q
P Q R
A Sales Price $213 $254 $200
B Raw material $35.00 $52.50 $17.50
C Direct labor $16.00 $12.00 $8.00
D=B+C Total direct cost $51.00 $64.50 $25.50
E=A-D Contribution per unit $162.00 $189.50 $174.50
F=E/A Contribution margin ratio 76.06% 74.61% 87.25%
2 Yes,Product R is the most profitable product
3 With this pricing Strategy, competitors willalsoincrease the price of R
and reduce price of P.
4 P Q R Total
A Annual Sales in units 8000 15000 4000
B Raw material Cost per unit $35.00 $52.50 $17.50
C=A*B Annual Raw Material Cost $280,000 $787,500 $70,000 $1,137,500
D=C/1137500 Proportion of total Raw mterial cost                 0.25           0.69            0.06
Allocation of Material Handling Cost as Percentage 25% 69% 6%
5 Allocation to P Allocation to Q Allocation to R
A B C=A*B D E=A*D F G=A*F
TotalCost % Amount
Machinery $1,225,000 25% $306,250 50% $612,500 25% $306,250
Machine set up $5,250 20% $1,050 30% $1,575 50% $2,625
Inspection $525,000 15% $78,750 45% $236,250 40% $210,000
Material Handling $875,000 25% $218,750 69% $603,750 6% $52,500
Engineering $344,750 35% $120,663 10% $34,475 55% $189,613
H Total $2,975,000 $725,463 $1,488,550 $760,988
I Sales in units          8,000           15,000          4,000
J=H/I Manufacturing Overhead per unit $90.68 $99.24 $190.25
K Total Direct Cost Per Unit $51.00 $64.50 $25.50
L=J+K Product Cost per unit based on Activity Based Costing $141.68 $163.74 $215.75
M=150%*L Budgeted Price $212.52 $245.61 $323.62